Let f be the function:


f(x) = ln(2x)-(2x2 +3), x > 0

(a) Use the sign pattern for f'(x) to determine the intervals where f rises and where f falls.


(b) Determine the coordinates of the local extreme point(s).


(c) Find f''(x) and determine where the graph of f is concave up and where it is concave down.


(d) Find any inflection points




1
Expert's answer
2021-08-03T07:13:39-0400

Let "f" be the function defined as "f(x) = \ln(2x)-(2x^2 +3),\ x > 0."


(a) The derivative of the function "f" is "f'(x) = \frac{2}{2x}-4x=\frac{1-4x^2}{x}=\frac{(1-2x)(1+2x)}{x}." Since "x>0," the equality "f'(x)=0" implies "x=\frac{1}{2}." If "x\in (0,\frac{1}{2})" then "f'(x)=\frac{(1-2x)(1+2x)}{x}>0." If "x>\frac{1}{2}," then "f'(x)<0." Therefore, on the interval "(0,\frac{1}{2})" the function "f" rises, on the interval "(\frac{1}{2},+\infty)" the function "f" falls.


(b) It follows from the previous item that "x=\frac{1}{2}" is a point of maximum, "f(\frac{1}{2})=\ln 1-(\frac{1}{2}+3)=-\frac{7}{2}."


(c) Let us find "f''(x)":

"f''(x)=(\frac{1}{x}-4x)'=-\frac{1}{x^2}-4=-\frac{1+4x^2}{x^2}."

It follows that "f''(x)<0" for all "x>0," and hence on the interval "(0,+\infty)" the graph of "f" is concave down.


(d) It follows from the previous item that the function "f" has no inflection point.